OpenStudy (anonymous):

. Are the graphs of 2y = -3x + 6 and y = 2/3 x + 2, parallel, Perpendicular, or neither? How do you know?

OpenStudy (anonymous):

You can tell by comparing the slopes.

OpenStudy (anonymous):

ik but idk how to graph them

OpenStudy (anonymous):

Put both in y=mx+b form

OpenStudy (anonymous):

The equations are perpendicular

OpenStudy (anonymous):

using the y=mx+b formula,m=-2/3 and b=3

OpenStudy (anonymous):

k then?

OpenStudy (anonymous):

The equations are perpendicular because the slopes of the two lines are negative reciprocals

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

What does the graph of x = -1 look like

OpenStudy (anonymous):

help me with this last one??

OpenStudy (anonymous):

@shortielynn , any equation of the form x=c is a vertical line through that number.

OpenStudy (anonymous):